15.6 g of an unknown molecular solute is dissolved in 5.02 moles of ethanol (C2H5OH), causing the freezing point of ethanol to decrease by 1.14 °C. Calculate the molar mass of the solute. For ethanol, the molar mass is 46.07 g/mol and Kf = 1.99 °C/m.
Answer #1

T = Kf × m

Kf = 1.99℃/m

m = molality = moles of solute / kg of solvent

We know moles = mass / molar mass

Mass of ethanol = 5.02 × 46.07 = 231.27 g = 0.2313 kg

Moles of solute = (∆T/Kf) × mass of solvent in kg = (1.14/1.99) × 0.2313 = 0.1325 mole

Molar mass of solute = mass in g / moles = 15.6 / 0.1325 = 117.74 g/mol

Thus answer: molar mass of solute = 117.74 g/mol